OpenStudy (anonymous):

MEDAL AND FAN FROM TWO ACCOUNTS PLEASE PLEASE PLEASE HELP Find all solutions to the equation. 7 sin2x - 14 sin x + 2 = -5

OpenStudy (mokeira):

let sinx=y \[7y ^{2}-14y+2=-5\] \[7y ^{2}-14y+7=0\] divide evrything by 7 to simplify \[y ^{2}-2y+1=0\] Factor \[(y-1)(y-1)=0\] y-1=0 y=1 sinx=y=1 sinx=1

OpenStudy (mokeira):

\[x=\sin ^{-1}1=90\]

OpenStudy (anonymous):

@Mokeira Thats the only answer? 90 deg?

OpenStudy (mokeira):

That is my answer